# Gravity and fluid dynamic correspondence on a null hypersurface: inconsistencies and advancement

### v1

@@ Abstract
Unlike black hole thermodynamics, the fluid-gravity correspondence of a generic null surface appears to be “incomplete”; though both approaches point to the emergent nature of gravitation. In the existing formulation of fluid-gravity correspondence of a null surface, we have only the momentum conservation relation in the form of the Damour-Navier-Stokes (DNS) equation. The other relations, such as the energy conservation relation, continuity relation, etc., are almost non-existent and are discussed sporadically in literature for stretched horizons. Furthermore, the fluid-gravity correspondence is not formulated from a suitable energy-momentum tensor in the conventional manner. In the paper, we address these issues. By giving the energy conservation relation, we improve the fluid-gravity analogy on a generic null hypersurface. Additionally, by introducing a suitable energy-momentum tensor, which forms the basis of the entire fluid description, we formulate the entire fluid description of the null surface. Moreover, we provide the expression of continuity relation in the context of fluid-gravity correspondence. The entire analysis has been done for a generic null surface, which differs from the existing stretched horizon approach and, thereby, expands the scope of applicability for any arbitrary null surface.
@@ 1 Introduction
Numerous signs in the literature point to the possibility that gravitation is not a fundamental force but rather an “emergent phenomenon”. The concept came from Sakharov’s work [1], and numerous other researchers later supported it [2, 3, 4]. These are some of the arguments that support this paradigm: (i) The thermodynamic laws can be directly linked to the field equations of many classes of gravitational theories [5, 6, 7, 8, 9, 10, 11, 12], (ii) The action functional of various theories can be identified as the free energy of the spacetime [13, 14, 15, 16, 17], (iii) Obtaining field equations from a thermodynamic extreme principle [18, 19], (iv) Possibility of obtaining the idea of energy equipartition theory for defining the microscopic degrees of freedom [20, 21, 22], (v) Dynamical equations for gravitation manifesting as the Navier-Stokes-like equation for wide class of theories [23, 24, 25, 26, 27, 28], (vi) Association of every solution of Navier-Stokes (NS) equation in $d$ dimension with a “dual” solution of Einstein’s equation in $d+1$ dimensions [29] (also see [30, 31, 32]) etc. By taking various projections of Einstein’s equation on the null surface, it is possible to establish the emergent nature of gravity more directly. Given that the normal, say $l^{a}$, generates the null surface, the auxiliary null vector is $k^{a}$, and the induced metric $q_{ab}=g_{ab}+l_{a}k_{b}+l_{b}k_{a}$ defines the null surface’s spatial cross-section, it can be demonstrated the following facts:
  • The projection $G_{ab}l^{a}l^{b}$ is related to Raychaudhuri’s equation, which plays a pivotal role in proving the second law of black hole (BH) thermodynamics (namely the area increase theorem).
  • For the metric in Gaussian null coordinate (GNC), the projection $G_{ab}l^{a}k^{b}$ was shown [10, 27] to manifest as the first law. Later, this has been established in a covariant manner [33] and also extended to scalar tensor theory [12] and Einstein-Cartan theory [34], thereby, interpreting $G_{ab}l^{a}k^{b}$ as a thermodynamic relation.
  • The other projection $G_{ab}l^{a}q^{b}_{c}$ provides Damour-Navier-Stokes (DNS) equation [23, 24], which is related to the hydrodynamic description of gravitational dynamics on the null surface. Such interpretation works beyond Einstein’s gravity (e.g. see [28, 35]).
All of the aforementioned phenomena demonstrate, in general, how gravitational theories are similar to two other completely uncorrelated important theories of (emergent) physics: fluid dynamics and thermodynamics. Notwithstanding several drawbacks, including the inability to find agreement regarding the microscopic degrees of freedom, the thermodynamic interpretation is remarkably reliable and consistent. This is not the case and even not fully complete, however, when considering fluid-gravity correspondence for a generic null surface and interpreting gravity as an emergent phenomenon. The DNS equation is currently the only relation that, from an emergent-gravity perspective, can support fluid-gravity correspondence (for a generic null surface). This DNS equation is related to the dynamical equation of a generic null surface and looks similar to the dynamical equation of fluid i.e. momentum conservation relation of the Navier-Stokes equation. Since both fluid, as well as gravity, still remain far from being completely well understood, one expects such a tantalizing connection might reveal several crucial information of about each other, i.e. acquiring knowledge in one domain might provide insight for a better understanding of another.  [^4] In fact, it is because of the earlier connection that drives us in this work to find analogous relations, which are there in fluid dynamics but are missing in null hydrodynamics. However, there appear several missing links and inconsistencies in the pursuit of correlating gravity with Navier-Stokes fluid. In the following, we enlist them.
  1. 1.
    As mentioned above, the main idea that revolves around the correspondence of gravitation with fluid dynamics is the fact that the gravitational equations (such as Einstein’s equation in GR) take the form of an NS-like equation when it is projected on a null surface and it is known as the DNS equation among the physicists. Although NS and DNS equations look alike, the NS equation and the DNS equation are not exactly equivalent. DNS equation contains a Lie-derivative as opposed to the convective derivative in the NS equation, and if the Lie-derivative is replaced by the convective derivative, there appear terms that do not have any fluid-dynamic interpretation. Although Padmanabhan [26] demonstrated that these extra terms, which undermine the fluid-dynamic argument, disappear in a local inertial frame when a particular metric is used, there is no justification for why this frame must be chosen.
  2. 2.
    As previously stated, the DNS equation resembles the fluid mechanics momentum conservation equation. The other crucial relations, like the “continuity equation” or the “energy conservation relation,” are not yet available, in contrast to classical fluid mechanics. Therefore, the fluid-gravity analogy’s emergent-gravity perspective (for a generic null surface) appears to be lacking. Hence, for a robust understanding and complete description, one must look for the existence of similar relations from the dynamics of gravity.
  3. 3.
    In the description of relativistic fluid dynamics, the energy-momentum tensor of fluid (EMT) is extremely crucial. The entire fluid dynamics revolve around the EMT. In fact, the energy conservation and the momentum conservation relation is obtained directly by taking a suitable projection of EMT conservation (i.e., $\nabla_{a}T^{ab}_{(fluid)}=0$). A similar energy-momentum tensor is also present in the membrane paradigm [36] and in Carroll dynamics [37, 38], both of which seeks to have a fluid interpretation for the stretched horizon, which is, generally, a timelike surface and becomes null at certain limit. On the contrary, in the context of a generic null surface, which was shown earlier (via DNS equation) to have a connection with NS fluid, a systematic approach using EM tensor is not yet introduced. Thus, one can ask whether a similar energy-momentum tensor exists for the hydrodynamic description of a generic null surface. If such an energy-momentum tensor exists, one can put null hydrodynamics and conventional fluid mechanics on equal footing. In addition, it can be considered to be a significant development as it can be applicable to any null surface (as opposed to the timelike surface being null only at a certain limit).
  4. 4.
    While correlating the DNS equation with the NS equation, the geometric quantities are identified with the fluid parameters (such as the non-affinity parameter the pressure, and so on). However, the consistencies of such identifications have not been checked thoroughly. One needs to look for other hydrodynamic equations of the null surface (other than the existing DNS equation) and check whether previous identifications are consistent with the new equations. In doing so, one can obtain a robust understanding of fluid-gravity correspondence.
  5. 5.
    Although DNS equation is consistent with the Kovtun-Son-Starinets (KSS) bound (i.e. $\eta/s\geq 1/4\pi$, where $\eta$ is the shear viscosity coefficient and $s$ is the entropy density), the bulk viscosity coefficient ($\zeta$) in DNS equation turns out to be negative (in GR, $\zeta=-1/16\pi$), implying the instability of fluid while subject to compression or expansion. Therefore, again one needs to find other hydrodynamical equations of null surfaces to check whether the earlier identification is correct.
Even if one accepts a specific choice of coordinate to relate DNS with NS (as proposed by Padmanabhan [26]), the other issues, especially 2 and 3, remain unaddressed. Until one resolves these issues, the fluid interpretation of a null hypersurface remains incomplete. This is because in three space dimensions the usual fluid is described by five independent parameters which are determined by five equations. The dynamical equation of fluid, which is the momentum conservation equation (namely NS), provides three equations; while the energy conservation and continuity relations give other two equations. Other than this, the NS fluid is the non-relativistic version of relativistic viscous fluid. The latter one is described through an energy-momentum tensor and correspondingly a non-relativistic version of this tensor exists for NS. Till now this direction has not been illuminated and, therefore, gravity-fluid connection, from the viewpoint of emergent phenomenon, cannot be considered to be on equal footing. Thus, the connection between gravity and fluid lacks a clear understanding and complete picture, unlike the thermodynamic description. For all these reasons, the fluid-gravity correspondence must be re-investigated thoroughly. Our primary objectives in this paper are twofold: (i) Finding an analogous of EMT, which should lie in the center while obtaining the hydrodynamical description of gravity, and (ii) Obtaining hydrodynamical equations on the null surface other than the DNS equation and check whether the earlier identifications are consistent.
With regard to finding EMT for a generic null hypersurface, the following background ideas can be helpful. It is known that for a spacelike surface $\mathcal{T}$ (defined by a timelike normal, say $n_{a}$), true dynamical degrees of freedom are the components of induced metric (not of the metric tensor) and the true dynamical information are embedded in its conjugate quantity. The situation is subtle for the null surface as one cannot define an induced metric on the null surface due to its degenerate property. However, accounting for an auxiliary null vector, one can define an induced metric $q_{ab}$ on the cross-section of the null surface (upon which both the null generator $l^{a}$ and the auxiliary null vector $k^{a}$ are the normals). Then, it can be argued that the true dynamical quantities for the null surface are $q_{ab}$ and the null generator $l^{a}$ (for more details, see [39, 40]). Therefore, the true dynamical information should be embedded in the quantities $\mathcal{P}_{ab}$ and $\mathcal{P}^{(l)}_{a}$, which are conjugate to $q^{ab}$ and $l^{a}$ respectively. Thus, we expect the possible hydrodynamical description on a null surface should be connected to both $\mathcal{P}_{ab}$ as well as $\mathcal{P}^{(l)}_{a}$. Recently, in literature [41, 42, 43] the Brown-York-like tensor is defined for the null surface which consists of both $\mathcal{P}_{ab}$ and $\mathcal{P}^{(l)}_{a}$. We find that such a gravitational quantity is a suitable entity that plays the analogous role of EMT for the effective fluid description of gravity on the null hypersurface.
Pertaining to the other objective, let us provide the following discussions. The other relations, that indicated fluid-gravity analogy, such as the energy conservation, continuity equation, etc. are discussed sporadically in literature. It is noteworthy that the previous attempts are formulated for the stretched horizon, which is null only at a certain limit. The formulations on the stretched surface have two distinct approaches. The old approach relates gravity with NS fluid via membrane paradigm [44, 45, 46] and the recent approach relates gravity with Carroll fluid [37, 38]. As mentioned earlier, these two formalisms are constructed for the stretched horizon and have not been generalized for a generic null surface. As a result, the existing approaches for the generic null surface merely rely on the similarity of the DNS equation with the momentum conservation relation of the NS equation and, hence, is incomplete (for the reasons mentioned earlier).
@@
In this paper, we provide a systematic description of fluid-gravity correspondence for a generic null surface. Our argument does not rely merely on finding the similarity between null dynamical equations with the Navier-Stokes equation and identifying the fluid parameters based on analogy. Instead, we follow a conventional approach of fluid dynamics, which is centered around the null Brown-York energy-momentum tensor– similar to the fact that the fluid-dynamic description is centered around the energy-momentum tensor of the fluid. We show that some of the fluid parameters can be identified naturally from the different components of null BY tensor, which agrees with the previous identification which was based on analogy. Furthermore, embracing the usual route of fluid dynamics, where the fluid equations (such as the energy conservation equation, momentum conservation equation, etc., which are known as the Navier-Stokes equation) can be obtained from the suitable projection of EMT, we show that the entire fluid-gravity description for a generic null surface can be formulated from the conservation of null BY tensor. Using this approach, not only do we obtain the age-old Damour-Navier-Stokes (DNS) equation, but we also show that the other relations, such as the energy conservation relation and the continuity equation can also be obtained.
@@
The paper is presented in the following parts. In the following 2, we have discussed the prerequisite knowledge on the part of fluid dynamics as well as on the geometry of the null surface, which will be useful for subsequent discussions. In 3, we introduce the Brown-York formalism– first the original one (valid for the timelike surface) and then its recent null extension. In 4, we discuss how the null Brown-York-like tensor provides the description of fluid-gravity correspondence, where we derive the Damour-Navier-Stokes equation, continuity equation as well as the Null-Raychaudhuri equation from the null Brown-York tensor. In 5, we show the connection of our result with the recent works on Carroll dynamics. Finally, we provide the discussions related to our analysis in 6.
@@
Let us describe the notations and conventions used. We will be working with the mostly positive (,+,+,+)(-,+,+,+)( - , + , + , + ) signature of the metric and employ the geometrized unit system where $c,\hbar$ and $G$ have been set to one. The lowercase Roman alphabets $a,b,\cdot\cdot\cdot$ represent the four-dimensional spacetime indices. The spatial indices on any three-dimensional spacelike surface are designated with the lowercase Latin alphabets $\mu,\nu,\cdot\cdot\cdot$. The indices on the null surface will be designated with the lowercase Latin alphabets, with a tilde over them $\tilde{\mu},\tilde{\nu},\cdot\cdot\cdot$. The uppercase Roman alphabets $A,B,\cdot\cdot\cdot$ will be reserved for the coordinates of a spatial two-dimensional cross-section/cut of the null surface.
@@ 2 Preambles and prerequisites
2.1 Relativistic fluid and its non-relativistic counterpart
In the discussion of (relativistic) fluid mechanics, the constitutive relations are obtained from an energy-momentum tensor (EMT) and baryon current. EMT encapsulates the density and flux of energy and momentum in spacetime, providing a comprehensive description of the distribution and flow of energy and momentum within a fluid. In addition to that, the energy-momentum tensor has the following properties: First, it is a conserved quantity i.e.,
$\nabla_{a}T^{ab}_{(fluid)}=0~{}.$ (1)

### v2

# A note on gravity and fluid dynamic correspondence on a null hypersurface

@@ Abstract
In the extensive literature on fluid-gravity correspondence formulated on null hypersurfaces, the Carrollian and membrane paradigm approaches have predominantly employed a timelike foliation. By contrast, within the null foliation formalism, only the momentum conservation law, expressed through the Damour-Navier-Stokes (DNS) equation, has been established. In this work, we revisit the null foliation formalism for a generic null hypersurface and extend it to include the energy conservation law, continuity equation, and related relations, all derived from the covariant conservation of an appropriately defined energy-momentum tensor. This development complements the existing literature on the fluid description of gravitational dynamics in the null foliation framework.
@@ 1 Introduction
Numerous signs in the literature point to the possibility that gravitation is not a fundamental force but rather an “emergent phenomenon”. The concept came from Sakharov’s work [1], and numerous other researchers later supported it [2, 3, 4]. These are some of the arguments that support this paradigm: (i) The thermodynamic laws can be directly linked to the field equations of many classes of gravitational theories [5, 6, 7, 8, 9, 10, 11, 12], (ii) The action functional of various theories can be identified as the free energy of the spacetime [13, 14, 15, 16, 17], (iii) Obtaining field equations from a thermodynamic extreme principle [18, 19], (iv) Possibility of obtaining the idea of energy equipartition theory for defining the microscopic degrees of freedom [20, 21, 22], (v) Dynamical equations for gravitation manifesting as the Navier-Stokes-like equation for wide class of theories [23, 24, 25, 26, 27, 28], (vi) Association of every solution of Navier-Stokes (NS) equation in $d$ dimension with a “dual” solution of Einstein’s equation in $d+1$ dimensions [29] (also see [30, 31, 32]) etc. All of the aforementioned phenomena broadly illustrate how gravitational theories bear striking similarities to two seemingly unrelated pillars of emergent physics: fluid dynamics and thermodynamics. Despite several limitations—most notably, the lack of consensus on the nature of the underlying microscopic degrees of freedom—the thermodynamic interpretation of gravity has proven to be remarkably robust and internally consistent. The same seems to be not so, however, for the fluid-gravity correspondence, especially when extended to generic null surfaces and the broader interpretation of gravity as an emergent phenomenon.
The connection between gravity and fluid dynamics was first uncovered by Damour [23, 24], whose work on null surfaces led to the membrane paradigm, treating black hole horizons as fluid-like systems [33]. This analogy matured with the AdS/CFT correspondence, where gravitational perturbations in AdS spacetimes were shown to map onto relativistic hydrodynamics on the boundary [34, 35, 36, 37], culminating in a precise correspondence between incompressible Navier–Stokes solutions and vacuum Einstein solutions [29]. Around the same time, Padmanabhan offered a thermodynamic perspective, demonstrating that Einstein’s equations on null surfaces can be derived from entropy extremization and are equivalent to the Navier–Stokes equations, thereby framing gravity as an emergent thermodynamic phenomenon that also reproduces key results such as the KSS bound [26].
A major breakthrough in this direction appeared when is was shown that the gravitational dynamics of a stretched surface, which is located arbitrarily close to an internal null boundary, say the event horizon, is similar to that of the Carrollian fluid [38, 39, 40, 41, 42, 43, 44]. A host of recent works [45, 46, 40, 47, 48, 49, 50, 51] speak in favour of this formalism, which is based on stretched horizon- a timelike proxy for the event horizon. Subsequent work in this gave rise to the Carrollian membrane paradigm. In relativistic fluid dynamics, the energy-momentum tensor (EMT) plays a central role, forming the foundation upon which the entire fluid description is constructed. The fundamental conservation laws of energy and momentum emerge naturally from appropriate projections of the EMT conservation equation $\nabla_{a}T^{ab}_{\rm(fluid)}=0$. A comparable EMT structure also appears in the membrane paradigm [33], and in recent formulations of Carrollian hydrodynamics for the stretched horizon [39, 43].
The gravitational dynamics of an internal null boundary at finite distances in the spacetime $(\mathcal{M},\textbf{g},\bm{\nabla})$ is naturally associated to fluid dynamics under different paradigms as mentioned above. At any point $p\in\mathcal{M}$ of spacetime, we can have a local approximate Rindler horizon corresponding to an accelerated observer. Such a Rindler horizon is generated by an approximate time-like boost Killing vector field whose norm vanishes on the Rindler horizon. Because the Rindler horizon is a one-way null surface, the Rindler observer can associate local thermodynamic entities associated with the event $p$ of the spacetime. A generic integrable null surface can locally be approximated by a Rindler horizon [10]. Hence, the dynamics of the null surface can locally be studied by such Rindler observers. The essential difference that arises in the paradigms mentioned above is due to the way the spacetime manifold in the neighborhood of $\mathcal{H}$ is foliated. In both the membrane paradigm and the Carroll membrane paradigm, the spacetime in the vicinity of the internal null boundary is foliated by a family of timelike hypersurfaces with the null limit leading to $\mathcal{H}$. However, following Carter [52], there can also be another foliation of $\mathcal{H}$ by the family of null hypersurfaces. The gravitational dynamics of $\mathcal{H}$ that need to be studied are independent of such a null foliation. In the Carrollian membrane paradigm, that exists in literature, the vicinity of the null boundary $\mathcal{H}$ (located at a finite distance) is foliated by a family of three-dimensional timelike surfaces $\mathcal{T}$. Such a family of timelike surfaces is defined by the scalar function $r(x^{a})=\text{constant}>0$. The null boundary is obtained in the limit $r(x^{a})=0$. That is, essentially, the null limit to $\mathcal{H}$ is obtained by taking the limit $r\rightarrow 0$. However, such a family $\mathcal{T}$ is not foliated by spacelike cross-sections/cuts (which would in turn provide the notion of a time evolution vector field $\bm{t}$ and hence predispose the structures $\mathcal{H}$ and $\mathcal{T}$ with a Galilean picture). Instead, Carroll structures are imposed on these timelike surfaces $\mathcal{T}$ as well as the null boundary $\mathcal{H}$ [39, 43]. With such a Carrollian structure implemented on $\mathcal{T}$, an appropriate Carrollian fluid energy-momentum tensor can be described on the stretched horizon [39, 43]. In the null limit, this Carrollian (surface) energy-momentum tensor matches with the one introduced in [53] up to an overall negative sign. The (vacuum) Einstein field equations are then shown to be conservation laws of such a Carrollian fluid energy-momentum tensor. The corresponding viscous stress tensor, pressure, energy density, momentum density, and heat current have been derived for such a Carrollian fluid [39, 43].
Consequently, a key open question remains: Does there exist a similar fluid description for the paradigm, where the spacetime in the neighborhood of the internal null boundary $\mathcal{H}$ is foliated by a family of null surfaces as opposed to timelike foliation adopted so far in the Carroll membrane paradigm? Given the significant difference in the construction of foliation, it is worthwhile to analyze the fluid description of the gravitational dynamics starting out with a possible fluid energy-momentum tensor as the usual starting point in the null foliation case. On the other hand within the analysis of null surfaces initiated by Damour and further developed by Padmanabhan, it has been shown that a momentum conservation equation—famously known as the Damour-Navier-Stokes (DNS) equation—can be derived from gravitational dynamics. However, unlike in classical fluid dynamics, other fundamental relations such as the continuity equation and the energy conservation law have not yet been established within this framework. Moreover, the whole classical fluid description is related to the conservation of fluid energy-momentum tensor. This gap indicates that the fluid-gravity correspondence, especially in its emergent gravity interpretation for generic null surfaces, remains incomplete. To achieve a more robust and comprehensive description, it is crucial to investigate whether these additional conservation laws can also be derived.
@@
In this paper, we provide a systematic description of fluid-gravity correspondence for a generic null surface. Our work proposes a novel extension of fluid-dynamic techniques that have recently been developed for the stretched horizon [33, 39, 43], adapting them to the unique characteristics of null surfaces. Unlike the stretched horizon, which retains a timelike nature, a null surface poses significant challenges due to its degenerate geometry and the absence of a well-defined rest frame. By addressing these issues, our approach provides a more comprehensive framework for describing the fluid-like behavior of null surfaces, broadening the scope of hydrodynamic descriptions in gravitational systems. Our argument does not rely merely on finding the similarity between null dynamical equations with the Navier-Stokes equation and identifying the fluid parameters based on analogy. Instead, we follow a conventional approach of fluid dynamics, which is centered around the null Brown-York energy-momentum tensor– similar to the fact that the fluid-dynamic description is centered around the energy-momentum tensor of the fluid. We show that some of the fluid parameters can be identified naturally from the different components of null BY tensor, which agrees with the previous identification which was based on analogy. Furthermore, embracing the usual route of fluid dynamics, where the fluid equations (such as the energy conservation equation, momentum conservation equation, etc., which are known as the Navier-Stokes equation) can be obtained from the suitable projection of EMT, we show that the entire fluid-gravity description for a generic null surface can be formulated from the conservation of null BY tensor. Using this approach, not only do we obtain the age-old Damour-Navier-Stokes (DNS) equation, but we also show that the other relations, such as the energy conservation relation and the continuity equation can also be obtained.
@@
The paper is presented in the following parts. In the following Section 2, we have discussed the prerequisite knowledge on the part of relativistic fluid dynamics as well as on the geometry of the null surface, which will be useful for subsequent discussions. In Section 3, we introduce the Brown-York formalism– first the original one (valid for the timelike surface) and then its recent null extension. In Section 4, we discuss how the null Brown-York-like tensor provides the description of fluid-gravity correspondence, where we derive the Damour-Navier-Stokes equation, continuity equation as well as the Null-Raychaudhuri equation from the null Brown-York tensor. In Section 5, we analyze the covariant gravitational dynamics w.r.t. an adapted coordinate system established on $\mathcal{H}$ and make connection with the non-relativistic fluid setting. Finally, we provide the discussions related to our analysis in Section 6.
@@
Let us describe the notations and conventions used. We will be working with the mostly positive (,+,+,+)(-,+,+,+) signature of the metric and employ the geometrized unit system where $c,\hbar$ and $G$ have been set to one. The lowercase Roman alphabets $a,b,\cdot\cdot\cdot$ represent the four-dimensional spacetime indices. The spatial indices on any three-dimensional spacelike surface are designated with the lowercase Latin alphabets $\mu,\nu,\cdot\cdot\cdot$. The indices on the null surface will be designated with the lowercase Latin alphabets, with a tilde over them $\tilde{\mu},\tilde{\nu}$$\cdot\cdot\cdot$. The uppercase Roman alphabets $A,B,\cdot\cdot\cdot$ will be reserved for the coordinates of a spatial two-dimensional cross-section/cut of the null surface. In Section 3, $\mathbf{u}_{a}$ represents arbitrary timelike/spacelike normal (in contrast to $s_{a}$, which is a spacelike normal). The tensor $\mathbf{h}_{ab}$ represents the induced metric on the corresponding arbitrary (timelike or spacelike) surface defined by $\mathbf{u}_{a}$ (whereas $h_{ab}$ is the induced metric of the timelike surface, described by the normal $s_{a}$). Similarly, the other boldfaced quantities, such as $\mathbf{K}_{ab}$, $\mathbf{T}^{a}$, $\boldsymbol{\pi}_{ab}$ etc., refer to geometric or physical quantities associated with this arbitrary timelike/spacelike surface described by $\mathbf{h}_{ab}$. In contrast, the non-bold quantities, such as $K_{ab}$, $T^{a}$, etc., are those associated specifically with the timelike surface whose induced metric is $h_{ab}$. Furthermore, boldface notation is sometimes used to indicate that the corresponding vector or tensor expressions hold irrespective of whether the indices are covariant or contravariant.
@@ 2 Preambles and prerequisites
2.1 Fluid equations